OpenStudy (anonymous):

5. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations

OpenStudy (anonymous):



OpenStudy (anonymous):

your slope is negative one , ends with +300 you have two poitnts already (0, 300) and (450, 0) y-y1= m(x-x1) y=-2/3 x + 300